# MC-NeRF: Multi-Camera Neural Radiance Fields for Multi-Camera Image Acquisition Systems

### v2

@@ 4.1   Dataset and Hardware Platfrom
TABLE 7: Quantitative camera parameters regression results of Fix-Step NeRF and MC-NeRF on all scenes.
In summary, the inclusion of linear transformations leads to both BARF and L2G-NeRF being unable to produce accurate results. Even in cases where calibration points are not provided, proposed method also faces challenges in achieving precision results. Notably, in the case of MC-NeRF without calibration points, the loss function comprises only the photometric loss and does not include the reprojection term.
4.6   Exp.5 Fix Step VS Global optimization
In this experiment, our goal is to validate the effectiveness of the global optimization framework and compare its performance against the Fixed Step NeRF rendering (Intrinsic and extrinsic parameters regression + NeRF, without other optimization). Fixed Step rendering simulates the typical steps involved in NeRF-based reconstruction using a multi-camera acquisition systems: Initially, camera intrinsic and extrinsic parameters are determined using calibration information. Subsequently, NeRF is trained using these fixed parameters to represent the 3D scene. Throughout the NeRF training phase, the camera parameters remain constant.
@@
In theory, global optimization is expected to provide more accurate intrinsic parameters, extrinsic parameters, and rendering results compared to the Fix Step method. In this experiment, both approaches utilized identical camera initialization parameters and employed the Adam optimizer with a learning rate of 0.01 for 20 epochs. This ensured that, upon entering the Global Optimization Stage in Fig.5, the camera parameters of the global optimization method closely resembled those of the Fix Step method.
The experimental results are presented in Fig.12, Table 7, and Table 7. We believe the global optimization has better performance compared to the Fix Step method. Regarding image rendering quality, in Fig.12, we randomly selected three scenes rendered in four different styles. The global optimization method stands out by producing sharper object boundaries and capturing finer object details. Table 7 shows the quantitative performance for 32 scenes. Firstly, MC-NeRF shows a significant performance improvement in Array Style, where it achieves the best scores across all three evaluation metrics. Secondly, we observed that MC-NeRF consistently outperforms in all LPIPS metrics. However, when it comes to PSNR and SSIM, except for the Array scene, MC-NeRF often has lower scores. We explored the reasons, as depicted in Fig.13.
Although MC-NeRF achieves better rendering details, it demonstrates larger errors in object boundaries alignment. We believe that the reason behind this phenomenon is the absence of enforced ray alignment constraints for rendering in MC-NeRF. Specifically, the camera parameters of Fix-Step NeRF are constant and remain stable during the rendering stage, which maintains a fixed ray distribution, and all fitting computations contained within NeRF’s MLPs. Based on the rendering results, it can be observed that to compensate for the ray distribution error resulting from inaccuracies in camera parameters, the rendered objects exhibit a multi-layer shadow structure. In Fig.13, the boundaries of the computer show noticeable shadows, while those of MC-NeRF appear much clearer and sharper. This multi-layer shadow structure has more advantages for pixel-to-pixel calculation methods like PSNR. If the clear-edged object generated by MC-NeRF is not aligned with the ground truth, the PSNR scores can be easily affected by the background, resulting in lower scores. Although MC-NeRF has some drawbacks in terms of boundary alignment, we prioritize high-quality rendering. Additionally, based on the rendering images, we believe that this alignment error is tolerable and unlikely to be readily noticeable.
Concerning the regression of camera parameters, as presented in Table 7, global optimization yields more accurate results with a significant improvement. We believe that during the global optimization, NeRF takes the lead as a supervisor for the regression. To achieve higher-quality rendering results, NeRF enforces the alignment of global rays, and this alignment constraint outperforms traditional camera calibration using calibration points. Additionally, the results also demonstrate that proposed decoupling constraint effectively resolves the coupling issue between intrinsic and extrinsic parameters in joint regression, allowing for the separation of parameters.
In summary, proposed method is proven to be effective. With global optimization, both the rendering images and the accuracy of regression parameters surpass those achieved using initial intrinsic and extrinsic parameters. Moreover, in this experiment, we noted that achieving higher-quality rendering results does not necessarily correlate with higher PSNR scores. We believe that this phenomenon is unique to the joint regression of intrinsic and extrinsic parameters. Because even when the regression results for both intrinsic and extrinsic parameters are highly accurate, their values are often challenging to align perfectly with ground truth. This leads to slight shifts of objects in the rendered image, which have a notable impact when calculating pixel-wise differences for PSNR. However, the LPIPS can still objectively reflect rendering performance even in such cases, as demonstrated by the conclusions in Table 7, where the results of global optimization consistently outperformed the Fix Step method.
5   CONCLUSIONS
In this paper, we present Multi-Camera Neural Radiance Fields (MC-NeRF) for Multi-Camera Image Acquisition Systems, a method that enables the joint optimization of both multi-camera intrinsic and extrinsic parameters during the training of neural radiance fields. We analyzed the challenges of joint optimization, including intrinsic parameters degenerated case and strong coupling between these parameters. Based on this analysis, we design the network architecture with training sequence, calibration objects, and a calibration data acquisition strategy. The proposed method primarily tackles the substantial workload associated with acquiring multi-camera intrinsic and extrinsic parameters when employing multi-camera acquisition systems for 3D scene representation. Code and models will be made available to the research community to facilitate reproducible research.
Although MC-NeRF allows for the regression of camera intrinsic and extrinsic parameters with NeRF, it still necessitates the acquisition of additional calibration data. We believe that the data collection scheme we propose can significantly reduce time requirements and enhance accuracy, but there is potential for optimization in the human-involved actions.
ACKNOWLEDGMENTS
This work is partly supported by National Natural Science Foundation of China (Grant No. 62233002, 61973034, U1913203 and CJSP Q2018229). The authors would like to thank Tianji Jiang, Xi Xu, Jiadong Tang, Zhaoxiang Liang, Xihan Wang, Dianyi Yang, and all other members of the ININ Lab of the Beijing Institute of Technology for their contribution to this work.

### v4

@@ ACKNOWLEDGMENTS
Figure 16: Rendering results of BARF+NeRF--- -, Fix-Step NeRF and MC-NeRF. The first row represents the ground truth images, while the remaining rows display rendering results. The rendering results indicate that simply stacking methods without designing an optimized sequence does not lead to improved performance.
TABLE 8: Quantitative rendering results of Mix-Step and MC-NeRF on all scenes. The comparison results show that our training sequence is effective, and simply stacking methods does not lead to better performance.
[…]
Mix-Step Computer 15.6289 14.1684 18.0732 22.0304 0.7937 0.8445 0.7438 0.9172 0.2376 0.2686 0.3228 0.0634
Ficus 17.4136 16.2707 21.0218 20.4416 0.9231 0.8959 0.9263 0.9479 0.1302 0.1752 0.1551 0.0941
Gate 19.0948 18.1299 23.5322 21.3187 0.9185 0.7781 0.9234 0.9358 0.1509 0.1874 0.1775 0.1045
Lego 20.7237 18.6574 21.0971 17.9275 0.8615 0.7870 0.7564 0.9099 0.1389 0.3503 0.2937 0.1658
Materials 18.6585 16.9294 24.4285 20.4265 0.9118 0.8764 0.9234 0.9531 0.1068 0.1386 0.0976 0.0753
Snowtruck 19.3878 20.0831 21.3667 16.7986 0.8891 0.9184 0.8259 0.9192 0.1417 0.0917 0.2615 0.1216
Statue 17.9503 19.1121 23.1452 17.5167 0.9231 0.8923 0.8876 0.9539 0.1161 0.1999 0.1767 0.0827
Train 18.6715 18.6559 18.8509 21.3719 0.8464 0.8810 0.9172 0.9175 0.1729 0.0966 0.2571 0.0875
Mean 18.4411 17.7508 21.4395 19.7290 0.8834 0.8592 0.8437 0.9318 0.1494 0.1885 0.2178 0.0994
Figure 17: Rendering results of COLMAP+NeRF, Meshroom+NeRF and MC-NeRF. The first row represents the ground truth images, while the remaining rows display rendering results. Based on the rendering results, SFM-based methods are not applicable across all scenes and use cases. In scenarios with sparse backgrounds or textures, the camera parameters obtained are often inaccurate or even fail entirely. In contrast, our method demonstrates greater generalization and reliability, though it does require additional image data, which increases the data collection workload.
TABLE 9: The Number of available images obtained by Different Methods. At the bottom of the table, $TotalNumber$ represents the total number of images in the current scene dataset, while the numbers at the top indicate the number of usable images after estimating camera poses using the respective methods. The table shows that in scenes with simple backgrounds or limited textures, SFM-based methods discard a large number of usable images, leading to reconstruction failures. In contrast, our method avoids this issue.
Scene COLMAP Meshroom AprilTag
Computer 2 2 2 100 3 37 26 99 110 100 88 100
Ficus 2 2 2 2 8 9 21 100 100 110 100 100
Gate 16 31 31 98 41 67 34 100 110 100 88 100
Lego 48 92 86 100 54 99 88 100 110 100 88 100
Materials 38 2 71 5 36 0 36 0 110 100 88 100
Snowtruck 33 96 28 100 63 97 84 100 110 100 88 100
Statue 2 20 31 100 14 56 65 100 110 100 88 100
Train 37 93 47 100 49 100 84 100 110 100 88 100
Total Number 110 100 88 100 110 100 88 100 110 100 88 100
TABLE 10: Comparison Rendering Results between Two SFM-based Composite Methods and MC-NeRF.
[…]
COLMAP+NeRF Computer - - - 18.5081 - - - 0.8906 - - - 0.1677
Ficus - - - - - - - - - - - -
Gate - 25.3121 21.1899 23.0474 - 0.9099 0.8889 0.9137 - 0.1206 0.2163 0.1195
Lego 20.8396 21.8068 22.2283 18.8225 0.9189 0.9074 0.8075 0.8749 0.1446 0.0494 0.1859 0.1936
Materials 22.5298 - 23.0754 - 0.8982 - 0.9577 - 0.1143 - 0.0957 -
Snowtruck 12.7945 26.4260 - 18.4196 0.7328 0.9155 - 0.8695 0.3671 0.0884 - 0.2072
Statue - - 22.1310 18.6054 - - 0.9243 0.8903 - - 0.1265 0.1205
Train 13.2663 27.0967 20.9158 18.4150 0.7214 0.9371 0.8816 0.8847 0.4036 0.0614 0.1205 0.1937
Mean 18.7213 25.1604 21.7981 19.4620 0.8499 0.9175 0.8920 0.8866 0.2087 0.0799 0.1490 0.1871
Meshroom+NeRF Computer - - - 21.5518 - - - 0.9081 - - - 0.0978
Ficus - - - 20.0217 - - - 0.9494 - - - 0.0990
Gate 17.8344 20.8989 22.219 21.3296 0.8807 0.8804 0.9110 0.9257 0.1852 0.2550 0.1642 0.2277
Lego 15.4833 18.3328 13.4810 22.2046 0.7851 0.7856 0.7227 0.9110 0.3341 0.2550 0.4191 0.1117
Materials 20.6267 - - - 0.9293 - - - 0.0998 - - -
Snowtruck 13.2374 16.9954 10.1618 19.4415 0.8998 0.8189 0.6761 0.9166 0.1758 0.2526 0.5043 0.1138
Statue - 22.9819 14.1917 - - 0.9684 0.8227 - - 0.0547 0.2509 -
Train 21.1296 17.8849 11.8588 21.8684 0.8911 0.7979 0.6913 0.9237 0.1323 0.2526 0.4561 0.0997
Mean 16.7955 19.4188 14.3824 21.0696 0.8737 0.8502 0.7647 0.9224 0.1987 0.1920 0.3589 0.1249
@@
Figure 19: Comparison of rendering performance in NeRF with mixed intrinsic dataset and single intrinsic dataset. The top row shows ground-truth images captured at varying FOVs, indicated below each image. The second row displays results produced by the NeRF model trained on each fixed FOV dataset, and the third row exhibits results from the NeRF model trained on a mixed dataset. The numbers in the top-right corner of each result represent the number of images in the dataset with the current FOV. In the mixed dataset, there are 22, 12, 10, 24, and 16 images for FOVs of 40, 50, 60, 70, and 80 degrees, respectively.
TABLE 11: Quantitative rendering results of SUPPLEMENTARY EXPERIMENT E.
$40^{\circ}$ $50^{\circ}$ $60^{\circ}$ $70^{\circ}$ $80^{\circ}$
[…]
Mixed FOV 24.169 27.333 29.949 32.138 33.786
Distance 1.271 0.539 0.053 0.797 0.997
[…]
Mixed FOV 0.819 0.899 0.939 0.963 0.975
Distance 0.039 0.014 0.003 0.006 0.005
[…]
Mixed FOV 0.147 0.069 0.037 0.026 0.016
Distance 0.039 0.013 0.003 0.009 0.002